For the cross elasticity I’d demand using arc method my book says the answer is -1.235 but I think it’s wrong . Can someone please help?? Am I wrong or is the book? What’s the cross elasticity of demand equation using arc method ?
1002 Introduction to economics Initially, the price of a tennis racket is £20. Demand is 30 and supply is 50. If the price falls by E5, the quanity demanded rises to 40, the quantity supplied rises to 40, and the quantity demanded of white cotton t-shirts rises from 70 to 100. Using the arc method calculate the own-price demand elasticity and the elasticity of supply for tennis racquets Activity SG3.9 cross-price demand elasticity for white cotton t-shirts.Are white cotton t-shirts a complement or substitute to tennis racquets? Own price elasticity = Cross-price elasticity Elasticity of supply Substitute or complement? BVFD: Examine Table 4.11 -this provides a simple but helpful summary of the chapter up to section 4.8. Make sure that you also understand section 4.7 which brings together own- price, cross-price and income elasticities with reference to inflation. idence of a tax BVFD: read section 4.9. The key point of this section is that the incidence of the tax is not to the person who physically pays the m whichever party (consumers or demand or supply) will bear the Suppose demand were perfectly tax be shared between consum It is important to realise that a sa paid by consumers (sometimes received by producers (the supply diagra in the a of taxes th An r is why g nt
Answer #1

After price fall, new price of tennis rackets (\pounds) = 20 - 5 = 15

Using Arc method,

Cross-price elasticity = (Change in quantity of T-shirts / Average quantity of T-shirts) / (Change in price of Tennis racket / Average price of Tennis rackets)

= [(100 - 70) / (100 + 70)] / [-\pounds5 / \pounds (20 + 15)]

= (30 / 170) / (- 5 / 35)

= (30 x 35) / [- (5 x 170)]

= 1,050 / (- 850)

= -1.235

Since cross-price elasticity is negative, T-shirts and Tennis rackets are complement goods.